OpenStudy (anonymous):

Help me solve this please.. -2/5+p=-4/15

OpenStudy (anonymous):

p = -4/15 + 2/5 = -4/15 +6/15 = 2 / 15 = 0.155555

OpenStudy (anonymous):

How'd you come up with that answer?

OpenStudy (anonymous):

The answer is a fraction

OpenStudy (anonymous):

2 / 15 then

OpenStudy (anonymous):

Ok hang on

OpenStudy (anonymous):

-2/5+p=-4/15 move the -2/5 to the right p = -4/15 + 2/5 make them both / 15 by making 2/5 = 6 / 15 p = - 4/15 + 6 / 15 p = 2 / 15

OpenStudy (anonymous):

Ok thanks.

OpenStudy (anonymous):

MattheyJ I have another equation, can you help me out iwth thisone?

OpenStudy (anonymous):

-6/7+z=-8/15